A glucose solution in water is labelled as 20%. The density of the solution is 1.20g/mL. What is the molarity if the solution?
1
Expert's answer
2020-04-03T11:51:51-0400

Mраствора=n(C6H12O6)/Vраствора


If the mass of the solution = 100g, then m(C6H12O6)=100*0.2=20g

n(C6H12O6)= m(C6H12O6)/M(C6H12O6)=20/180=0.11mol


m=ρ*V

V=m/ρ=100/1.2=83.3ml=0.083l


Mраствора=n(C6H12O6)/Vраствора=0.11/0.083=1.33mol/l